Identify the perfect squares among the following numbers
1, 2, 3, 8, 36, 49, 65, 67, 71, 81, 169, 625, 125, 900, 100, 1000, 10000

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Perfect square – It is the product of some integer with itself. These numbers are non-negative.For example, 9 is a square number, since it can be written as 3×3 Complete step-by-step answer: Many mathematical operations have an inverse, or opposite, operation. Subtraction is the opposite of addition, division is the inverse of multiplication, and so on. Squaring, which we learned about in a previous lesson (exponents), has an inverse too, called "finding the square root." Remember, the square of a number is that number times itself. The perfect squares are the squares of the whole numbers: 1, 4, 9, 16, 25, 36, 49, 64, 81, 100 … 12=1×1=1 62=6×6=36 72−7×7=49 92=9×9=81 132=13×13=169 252=25×25=625 302=30×30=900 102=10×10=100 So, from above given numbers, 1, 36, 49, 81, 169, 625, 900 and 100 are perfect squares. Note: In base 10, a square number can end only with digits 0, 1, 4, 5, 6 or 9 as follows : If the last digit of a number is 0, its square ends in 0 (in fact, the last two digits must be 00) If the last digits of a number is 1 or 9, its squares ends in 1 If the last digits of a number is 2 or 8, its squares ends in 4 If the last digits of a number is 3 or 7, its squares ends in 9 If the last digits of a number is 4 or 6, its squares ends in 6 If the last digits of a number is 5, its square ends in 5 (in fact, the last two digits must be 25).